OpenStudy (anonymous):

hey what is 5-2y, if y=-3

OpenStudy (anonymous):

1 cuz 2 x 3 is 6 6-5=1

TheSmartOne (thesmartone):

Just plug it in. xD

sammixboo (sammixboo):

Wel plug in the info 5 - 2y 5 - 2*-3 Following the order of operations, you multiply/divide first, but we only have multiplication, so what is 2*-3?

TheSmartOne (thesmartone):

Also \(\Huge\cal\color{red}{Welcome~to~Open~Study!!}\)

TheSmartOne (thesmartone):

@Anon101 You have been on OS for a while. Don't you know it is against the CoC to give direct answers. http://openstudy.com/code-of-conduct

sammixboo (sammixboo):

You are not doing 2*3, @Anon101 it is 2*-3 :)

OpenStudy (anonymous):

I explained it ok gosh...

sammixboo (sammixboo):

So it is not 1 :(

TheSmartOne (thesmartone):

\(\ 5-2y\) \(\ 5-2(-3)\)

sammixboo (sammixboo):

And plus if it was 6 you would have 5 - 6 not 6 - 5 and the answer would be -1 :)

TheSmartOne (thesmartone):

@lilbradley14

TheSmartOne (thesmartone):

\(\ - \times -= +\)

TheSmartOne (thesmartone):

So \(\ -2 \times -3 = +?\)